# Joint Transmission and Compression Design for 6G Networked Sensing with Limited-Capacity Backhaul

### v1

@@ III Posterior Cramér-Rao Bound for Localization
$\bm{U}\triangleq\frac{\partial\boldsymbol{\zeta}}{\partial\boldsymbol{\xi}}=\begin{bmatrix}\hat{\bm{U}}&\bm{0}\\ \bm{0}&\bm{I}\end{bmatrix}\in\mathbb{C}^{(2K+N^{2}K)\times(NK+N^{2}K)},$ (13)
@@
with $\bm{U}_{n}$ being a block diagonal matrix whose $k$-th block diagonal component is $\bm{u}_{n,k}$, $\forall n,k$. Moreover, $\bm{F}_{0,\boldsymbol{\zeta}}$ is defined as
$\bm{F}_{0,\boldsymbol{\zeta}}=\mathbb{E}_{\tilde{\bm{y}}}\left[\frac{\partial\log p(\tilde{\bm{y}}|\boldsymbol{\zeta})}{\partial\boldsymbol{\zeta}}\Bigg{(}\frac{\partial\log p(\tilde{\bm{y}}|\boldsymbol{\zeta})}{\partial\boldsymbol{\zeta}}\Bigg{)}^{T}\right],$ (15)
where $p(\tilde{\bm{y}}|\boldsymbol{\zeta})$ has the similar form to $p(\tilde{\bm{y}}|\boldsymbol{\xi})$, but we should treat $\bm{G}_{n,u}$’s in (2) as functions of $\bm{\theta}$ instead of $\bm{q}$. The explicit expression for $\bm{F}_{0,\boldsymbol{\zeta}}$ can be given by the following proposition.
@@
Define $\bm{F}_{1}$, $\bm{F}_{2}$, and $\bm{F}_{3}$ as (16) and (17) on the top of the next page, where
$\bm{F}_{1}$ $=\begin{bmatrix}\bm{F}_{1}(1,1)&\dots&\bm{F}_{1}(1,N)\\ \vdots&\ddots&\vdots\\ \bm{F}_{1}(N,1)&\dots&\bm{F}_{1}(N,N)\end{bmatrix},\bm{F}_{2}=\begin{bmatrix}\bm{F}_{2}(1,1,1)&\dots&\bm{F}_{2}(1,1,N)&\dots&\bm{F}_{2}(1,N,1)&\dots&\bm{F}_{2}(1,N,N)\\ \vdots&\ddots&\vdots&\dots&\vdots&\ddots&\vdots\\ \bm{F}_{2}(N,1,1)&\dots&\bm{F}_{2}(N,1,N)&\dots&\bm{F}_{2}(N,N,1)&\dots&\bm{F}_{2}(N,N,N)\end{bmatrix},$ (16)
$\bm{F}_{3}$ $=\begin{bmatrix}\bm{F}_{3}(1,1,1,1)&\dots&\bm{F}_{3}(1,1,1,N)&\dots&\bm{F}_{3}(1,1,N,1)&\dots&\bm{F}_{3}(1,1,N,N)\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \bm{F}_{3}(1,N,1,1)&\dots&\bm{F}_{3}(1,N,1,N)&\dots&\bm{F}_{3}(1,N,N,1)&\dots&\bm{F}_{3}(1,N,N,N)\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \bm{F}_{3}(N,1,1,1)&\dots&\bm{F}_{3}(N,1,1,N)&\dots&\bm{F}_{3}(N,1,N,1)&\dots&\bm{F}_{3}(N,1,N,N)\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \bm{F}_{3}(N,N,1,1)&\dots&\bm{F}_{3}(N,N,1,N)&\dots&\bm{F}_{3}(N,N,N,1)&\dots&\bm{F}_{3}(N,N,N,N)\end{bmatrix}.$ (17)
$\bm{F}_{1}(i,i)=\sum_{u=1}^{N}\Big{[}\Big{(}\bm{A}_{u}^{H}\bm{O}^{-1}_{u}\bm{A}_{u})\odot(\bm{B}^{*}_{ui}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\dot{\bm{V}}_{i}\bm{B}_{ui}\Big{)}\Big{]}$
$+(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}^{*}_{ii}\bm{V}_{i}^{H}\bm{R}_{i}^{*}\dot{\bm{V}}_{i}\bm{B}_{ii})$
$+(\bm{A}_{i}^{H}\bm{O}^{-1}_{i}\dot{\bm{A}}_{i})\odot(\bm{B}^{*}_{ii}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}\bm{B}_{ii})$
$+\sum_{u=1}^{N}\Big{[}\Big{(}\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\dot{\bm{A}}_{i})\odot(\bm{B}_{iu}^{*}\bm{V}_{u}^{H}\bm{R}_{u}^{*}\bm{V}_{u}\bm{B}_{iu}\Big{)}\Big{]},\forall i,$ (18)
$\bm{F}_{1}(i,n)=(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}^{*}_{in}\bm{V}_{n}^{H}\bm{R}_{n}^{*}\dot{\bm{V}}_{n}\bm{B}_{in})$
$+(\bm{A}_{n}^{H}\bm{O}^{-1}_{n}\dot{\bm{A}}_{n})\odot(\bm{B}^{*}_{ni}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}\bm{B}_{ni}),\forall i\neq n,$ (19)
$\bm{F}_{2}(i,i,i)=(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}_{ii}^{*}\bm{V}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i})$
$+(\bm{A}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}_{ii}^{*}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}),\forall i,$ (20)
$\bm{F}_{2}(i,i,n)=(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}_{in}^{*}\bm{V}_{n}^{H}\bm{R}_{n}^{*}\bm{V}_{n}),\forall i\neq n,$ (21)
$\bm{F}_{2}(i,n,i)=(\bm{A}_{n}^{H}\bm{O}^{-1}_{n}\bm{A}_{n})\odot(\bm{B}_{ni}^{*}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}),\forall i\neq n,$ (22)
$\bm{F}_{2}(i,n,k)=\mathbf{0},\forall i\neq n,i\neq k,$ (23)
$\bm{F}_{3}(n,i,n,i)=(\bm{A}_{n}^{H}\bm{O}^{-1}_{n}\bm{A}_{n})\odot(\bm{V}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}),\forall i,$ (24)
$\bm{F}_{3}(n,i,n,k)=\mathbf{0},\forall i\neq k,$ (25)
$\bm{F}_{3}(n,i,u,k)=\mathbf{0},\forall n\neq u,$ (26)
@@
with $\dot{\bm{A}}_{n}=[\frac{\partial\bm{a}(\theta_{n,1})}{\partial\theta_{n,1}},\ldots,\frac{\partial\bm{a}(\theta_{n,K})}{\partial\theta_{n,K}}]\in\mathbb{C}^{M_{r}\times K}$ and $\bm{O}_{n}=\sigma^{2}\bm{I}+\bm{Q}_{n}\in\mathbb{C}^{M_{r}\times M_{r}}$, $\forall n$. Then, $\bm{F}_{0,\boldsymbol{\zeta}}$ given in (15) can be expressed as
$\bm{F}_{0,\boldsymbol{\zeta}}=2\begin{bmatrix}\Re\{\bm{F}_{1}\}&\Re\{\bm{F}_{2}\}&-\Im\{\bm{F}_{2}\}\\ \Re\{\bm{F}_{2}\}^{T}&\Re\{\bm{F}_{3}\}&-\Im\{\bm{F}_{3}\}\\ -\Im\{\bm{F}_{2}\}^{T}&-\Im\{\bm{F}_{3}\}^{T}&\Re\{\bm{F}_{3}\}\end{bmatrix},$ (27)
@@
The details of proof are omitted due to limited pages. ∎
After characterizing the FIM $\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ shown in (5), where $\bm{F}_{0,\boldsymbol{\xi}}$ is obtained via Proposition 1 and $\bm{F}_{p}$ is obtained via (III), the PCRB for the mean-squared error (MSE) of any estimator of $\bm{q}$, denoted by $\hat{\bm{q}}$, can be expressed as
$\mathbb{E}_{\bm{q}}[\|\hat{\bm{q}}-\bm{q}\|^{2}]\geq\sum_{i=1}^{2K}[(\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N}))^{-1}](i,i),$ (28)
where for any squared matrix $\bm{X}$, $\bm{X}(i,i)$ denotes its $i$-th diagonal element.
@@
We want to emphasize that our result is a generalization of the result in [16]. Specifically, the work [16] has characterized the Cramér-Rao Bound for estimating the unknown but deterministic locations of targets for the special case of one BS, i.e., $N=1$, while our result holds for estimating the random target locations in the general case of networked sensing via cooperation among multiple BSs.
@@ IV Optimizing Transmission and Compression Strategies for PCRB Minimization
In this section, we aim to design the transmit covariance matrices and the compression noise covariance matrices of all the BSs, i.e., $\bm{R}_{n}$ and $\bm{Q}_{n}$, $\forall n$, to minimize the PCRB for localization, subject to the individual transmit power constraint and backhaul capacity constraint of each BS. Specifically, the above problem can be formulated as
@@
where $\bar{P}_{n}$ is transmit power constraint for BS $n$, $\bar{D}_{n}$ is the backhaul capacity constraint for BS $n$, and $D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bm{Q}_{n})$ is the backhaul transmission rate of BS $n$ given in (II), $\forall n$.
Note that the objective function of Problem (P1) is a complicated function of $\bm{R}_{n}$’s and $\bm{Q}_{n}$’s. To tackle this challenge, we adopt the Schur complement technique to equivalently transform Problem (P1) to the following problem
@@
s.t. $~{}\begin{bmatrix}\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})&\bm{e}_{i}\\ \bm{e}_{i}^{T}&t_{i}\end{bmatrix}\succeq\bm{0},~{}\forall i,$ (30b)
@@
where $t_{1},\ldots,t_{2K}$ are auxiliary variables and $\bm{e}_{i}$ is the $i$th column of the identity matrix $\bm{I}$, $\forall i$. Problem (P2) is still challenging, because all the variables are coupled together. In this work, we design an alternating optimization based algorithm to iteratively solve Problem (P2). Specifically, at each iteration of the alternating optimization based algorithm, we first fix $\bm{Q}_{n}$’s and optimize $\bm{R}_{n}$’s and $t_{i}$’s in Problem (P2), and then fix $\bm{R}_{n}$’s and optimize $\bm{Q}_{n}$’s and $t_{i}$’s in Problem (P2). This can guarantee that the objective value of Problem (P2) will be decreased after each iteration. The algorithm will terminate at some iteration when the decrease in the objective value is smaller than some threshold. In the following, we respectively show how to optimize $\bm{R}_{n}$’s and $t_{i}$’s given $\bm{Q}_{n}$’s and how to optimize $\bm{Q}_{n}$’s and $t_{i}$’s given $\bm{R}_{n}$’s in Problem (P2).
IV-A Optimization of Transmit Covariance Matrices

### v3

@@ III Posterior Cramér-Rao Bound for Localization
$\bm{U}\triangleq\frac{\partial\bm{\zeta}}{\partial\bm{\xi}}$ $=\begin{bmatrix}\hat{\bm{U}}&\bm{0}\\ \bm{0}&\bm{I}\end{bmatrix}\in\mathbb{C}^{(2K+N^{2}K)\times(NK+N^{2}K)},$ (14)
@@
$\bm{U}_{n}$ $=\textbf{blkdiag}\left\{\bm{u}_{n,1},\dots,\bm{u}_{n,K}\right\}$
$=\begin{bmatrix}\bm{u}_{n,1}&\dots&\bm{0}\\ \vdots&\ddots&\vdots\\ \bm{0}&\dots&\bm{u}_{n,K}\end{bmatrix}\in\mathbb{C}^{2K\times K},\forall n.$ (16)
Moreover, $\bm{F}_{0,\bm{\zeta}}$ is defined as
$\bm{F}_{0,\bm{\zeta}}=\mathbb{E}_{\tilde{\bm{y}}}\left[\frac{\partial\log p(\tilde{\bm{y}}|\bm{\zeta})}{\partial\bm{\zeta}}\Bigg{(}\frac{\partial\log p(\tilde{\bm{y}}|\bm{\zeta})}{\partial\bm{\zeta}}\Bigg{)}^{T}\right],$ (17)
where $p(\tilde{\bm{y}}|\bm{\zeta})$ has the similar form to $p(\tilde{\bm{y}}|\bm{\xi})$, but we should treat $\bm{G}_{n,u}$’s given in (3) as functions of $\bm{\theta}$ instead of $\bm{q}$.
The explicit expression for $\bm{F}_{0,\bm{\zeta}}$ can be given by the following proposition.
@@
Define $\bm{F}_{1}$, $\bm{F}_{2}$, and $\bm{F}_{3}$ as (18) – (20) on the top of the next page, where
$\bm{F}_{1}$ $=\begin{bmatrix}\bm{F}_{1}[1,1]&\dots&\bm{F}_{1}[1,N]\\ \vdots&\ddots&\vdots\\ \bm{F}_{1}[N,1]&\dots&\bm{F}_{1}[N,N]\end{bmatrix}\in\mathbb{C}^{NK\times NK},$ (18)
$\bm{F}_{2}$ $=\begin{bmatrix}\bm{F}_{2}[1,1,1]&\dots&\bm{F}_{2}[1,1,N]&\dots&\bm{F}_{2}[1,N,1]&\dots&\bm{F}_{2}[1,N,N]\\ \vdots&\ddots&\vdots&\dots&\vdots&\ddots&\vdots\\ \bm{F}_{2}[N,1,1]&\dots&\bm{F}_{2}[N,1,N]&\dots&\bm{F}_{2}[N,N,1]&\dots&\bm{F}_{2}[N,N,N]\end{bmatrix}\in\mathbb{C}^{NK\times N^{2}K},$ (19)
$\bm{F}_{3}$ $=\begin{bmatrix}\bm{F}_{3}[1,1,1,1]&\dots&\bm{F}_{3}[1,1,1,N]&\dots&\bm{F}_{3}[1,1,N,1]&\dots&\bm{F}_{3}[1,1,N,N]\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \bm{F}_{3}[1,N,1,1]&\dots&\bm{F}_{3}[1,N,1,N]&\dots&\bm{F}_{3}[1,N,N,1]&\dots&\bm{F}_{3}[1,N,N,N]\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \bm{F}_{3}[N,1,1,1]&\dots&\bm{F}_{3}[N,1,1,N]&\dots&\bm{F}_{3}[N,1,N,1]&\dots&\bm{F}_{3}[N,1,N,N]\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ \bm{F}_{3}[N,N,1,1]&\dots&\bm{F}_{3}[N,N,1,N]&\dots&\bm{F}_{3}[N,N,N,1]&\dots&\bm{F}_{3}[N,N,N,N]\end{bmatrix}\in\mathbb{C}^{N^{2}K\times N^{2}K}.$ (20)
$\bm{F}_{1}[i,i]=\sum_{u=1}^{N}\Big{[}\Big{(}\bm{A}_{u}^{H}\bm{O}^{-1}_{u}\bm{A}_{u})\odot(\bm{B}^{*}_{u,i}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\dot{\bm{V}}_{i}\bm{B}_{u,i}\Big{)}\Big{]}$
$+\sum_{u=1}^{N}\Big{[}\Big{(}\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\dot{\bm{A}}_{i})\odot(\bm{B}_{i,u}^{*}\bm{V}_{u}^{H}\bm{R}_{u}^{*}\bm{V}_{u}\bm{B}_{i,u}\Big{)}\Big{]}$
$+(\bm{A}_{i}^{H}\bm{O}^{-1}_{i}\dot{\bm{A}}_{i})\odot(\bm{B}^{*}_{i,i}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}\bm{B}_{i,i})$
$+(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}^{*}_{i,i}\bm{V}_{i}^{H}\bm{R}_{i}^{*}\dot{\bm{V}}_{i}\bm{B}_{i,i}),~{}\forall i,$ (21)
$\bm{F}_{1}[i,n]=(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}^{*}_{i,n}\bm{V}_{n}^{H}\bm{R}_{n}^{*}\dot{\bm{V}}_{n}\bm{B}_{i,n})$
$+(\bm{A}_{n}^{H}\bm{O}^{-1}_{n}\dot{\bm{A}}_{n})\odot(\bm{B}^{*}_{n,i}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}\bm{B}_{n,i}),~{}\forall i\neq n,$ (22)
$\bm{F}_{2}[i,i,i]=(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}_{i,i}^{*}\bm{V}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i})$
$+(\bm{A}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}_{i,i}^{*}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}),~{}\forall i,$ (23)
$\bm{F}_{2}[i,i,n]=(\dot{\bm{A}}_{i}^{H}\bm{O}^{-1}_{i}\bm{A}_{i})\odot(\bm{B}_{i,n}^{*}\bm{V}_{n}^{H}\bm{R}_{n}^{*}\bm{V}_{n}),~{}\forall i\neq n,$ (24)
$\bm{F}_{2}[i,n,i]=(\bm{A}_{n}^{H}\bm{O}^{-1}_{n}\bm{A}_{n})\odot(\bm{B}_{n,i}^{*}\dot{\bm{V}}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}),~{}\forall i\neq n,$ (25)
$\bm{F}_{2}[i,n,k]=\mathbf{0},~{}\forall i\neq n,i\neq k,$ (26)
$\bm{F}_{3}[n,i,n,i]=(\bm{A}_{n}^{H}\bm{O}^{-1}_{n}\bm{A}_{n})\odot(\bm{V}_{i}^{H}\bm{R}_{i}^{*}\bm{V}_{i}),~{}\forall i,$ (27)
$\bm{F}_{3}[n,i,n,k]=\mathbf{0},~{}\forall i\neq k,$ (28)
$\bm{F}_{3}[n,i,u,k]=\mathbf{0},~{}\forall n\neq u,$ (29)
@@
with $\dot{\bm{A}}_{n}=\left[\frac{\partial\bm{a}(\theta_{n,1})}{\partial\theta_{n,1}},\ldots,\frac{\partial\bm{a}(\theta_{n,K})}{\partial\theta_{n,K}}\right]\in\mathbb{C}^{M_{r}\times K}$, $\dot{\bm{V}}_{n}=\left[\frac{\partial\bm{v}(\theta_{n,1})}{\partial\theta_{n,1}},\ldots,\frac{\partial\bm{v}(\theta_{n,K})}{\partial\theta_{n,K}}\right]\in\mathbb{C}^{M_{t}\times K}$, and $\bm{O}_{n}=\sigma^{2}\bm{I}+\bm{Q}_{n}\in\mathbb{C}^{M_{r}\times M_{r}}$, $\forall n$. Then, $\bm{F}_{0,\bm{\zeta}}$ given in (17) can be expressed as
$\bm{F}_{0,\bm{\zeta}}=2\begin{bmatrix}\Re\{\bm{F}_{1}\}&\Re\{\bm{F}_{2}\}&-\Im\{\bm{F}_{2}\}\\ \Re\{\bm{F}_{2}\}^{T}&\Re\{\bm{F}_{3}\}&-\Im\{\bm{F}_{3}\}\\ -\Im\{\bm{F}_{2}\}^{T}&-\Im\{\bm{F}_{3}\}^{T}&\Re\{\bm{F}_{3}\}\end{bmatrix},$ (30)
@@
Please refer to Appendix A. ∎
After characterizing the PFIM $\bm{F}_{\bm{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ shown in (6), where $\bm{F}_{0,\bm{\xi}}$ is obtained via Proposition 1 and $\bm{F}_{p}$ is obtained via (III), the PCRB for the mean-squared error (MSE) of any estimator of $\bm{q}$, denoted by $\hat{\bm{q}}$, can be expressed as
$\mathbb{E}_{\bm{q}}[\|\hat{\bm{q}}-\bm{q}\|^{2}]\geq\sum_{i=1}^{2K}[(\bm{F}_{\bm{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N}))^{-1}](i,i).$ (31)
@@
We want to emphasize that our result is a generalization of the result in [13]. Specifically, the work [13] has characterized the CRB for estimating the unknown but deterministic locations of targets for the special case of one BS, i.e., $N=1$, while our result holds for estimating the random target locations in the general case of networked sensing via cooperation among multiple BSs.
@@ IV Optimizing Transmission and Compression Strategies for PCRB Minimization
In this section, we aim to design the transmit covariance matrices and the compression noise covariance matrices of all the BSs, i.e., $\bm{R}_{n}$ and $\bm{Q}_{n}$, $\forall n$, to minimize the PCRB for localization, subject to the individual transmit power constraint and fronthaul capacity constraint of each BS. Specifically, the above problem can be formulated as
@@
where $\bar{P}_{n}$ is transmit power constraint for BS $n$, $\bar{D}_{n}$ is the fronthaul capacity constraint for BS $n$, and $D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bm{Q}_{n})$ is the fronthaul transmission rate of BS $n$ given in (II), $\forall n$.
Our aim is to design the BSs’ transmit covariance matrices, i.e., $\bm{R}_{n}$’s, and compression noise covariance matrices, i.e., $\bm{Q}_{n}$’s, in Problem (P1). Note that the objective function of Problem (P1) is a complicated function of $\bm{R}_{n}$’s and $\bm{Q}_{n}$’s. To tackle this challenge, we adopt the Schur complement technique to equivalently transform Problem (P1) to the following problem
@@
s.t. $~{}\begin{bmatrix}\bm{F}_{\bm{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})&\bm{e}_{i}\\ \bm{e}_{i}^{T}&t_{i}\end{bmatrix}\succeq\bm{0},~{}\forall i,$ (33b)
@@
where $t_{1},\ldots,t_{2K}$ are auxiliary variables and $\bm{e}_{i}$ is the $i$th column of the identity matrix $\bm{I}$, $\forall i$. Problem (P2) is still challenging, because all the variables are coupled together. In this work, we design an AO-based algorithm to iteratively solve Problem (P2). Specifically, at each iteration of the AO-based algorithm, we first fix $\bm{Q}_{n}$’s and optimize $\bm{R}_{n}$’s and $t_{i}$’s in Problem (P2), and then fix $\bm{R}_{n}$’s and optimize $\bm{Q}_{n}$’s and $t_{i}$’s in Problem (P2). This can guarantee that the objective value of Problem (P2) will be decreased after each iteration. The algorithm will terminate at some iteration when the decrease in the objective value is smaller than some threshold. In the following, we respectively show how to optimize $\bm{R}_{n}$’s and $t_{i}$’s given $\bm{Q}_{n}$’s and how to optimize $\bm{Q}_{n}$’s and $t_{i}$’s given $\bm{R}_{n}$’s in Problem (P2).
IV-A Optimization of Transmission Covariance Matrices